OpenStudy (ny,ny):

Write an equation of a polynomial function, which satisfies the given conditions. (Draw a graph.) Degree = 3. Zeros are -5, 0, 5. Passes through (2,21)

OpenStudy (tkhunny):

Try this one... Write an equation of a polynomial function, which satisfies the given conditions. Degree = 1. Zero is 5. Passes through (2,21)

OpenStudy (ny,ny):

y2-y1/x2-x1 21-0/2-5 21/-3 -7 0 = -7(5) + b 0 = -35 + b 35 = b y = -7x+35 @tkhunny

OpenStudy (anonymous):

y = a(x - p)(x - q)(x - r) , where p, q, and r are the zeros to get \[y = ax(x + 5)(x - 5)\] Plug in the point to solve for \(a\)

OpenStudy (ny,ny):

Which point? @peachpi

OpenStudy (anonymous):

(2,21)

OpenStudy (ny,ny):

21 = a (2)(2+5)(2-5) 21 = 2a (7) (-3) 21 = 2a (-21) 21 = -42a 21/-42 = a a=-.5

OpenStudy (anonymous):

yes that's right, so just replace a with -0.5 to complete your function

OpenStudy (ny,ny):

y=-.5x(x+5)(x−5)

OpenStudy (anonymous):

yep

OpenStudy (ny,ny):

I have a question.

OpenStudy (anonymous):

what is it?

OpenStudy (ny,ny):

Nevermind, I understand. :p Thank you.

OpenStudy (anonymous):

ok. you're welcome